# The Coverage Principle: A Framework for Understanding Compositional Generalization

### v1

@@ Abstract
Large language models excel at pattern matching, yet often fall short in systematic compositional generalization. We propose the coverage principle: a data-centric framework showing that models relying primarily on pattern matching for compositional tasks cannot reliably generalize beyond substituting fragments that yield identical results when used in the same contexts. We demonstrate that this framework has a strong predictive power for the generalization capabilities of Transformers. First, we derive and empirically confirm that the training data required for two-hop generalization grows at least quadratically with the token set size, and the training data efficiency does not improve with 20x parameter scaling. Second, for compositional tasks with path ambiguity where one variable affects the output through multiple computational paths, we show that Transformers learn context-dependent state representations that undermine both performance and interpretability. Third, Chain-of-Thought supervision improves training data efficiency for multi-hop tasks but still struggles with path ambiguity. Finally, we outline a mechanism-based taxonomy that distinguishes three ways neural networks can generalize: structure-based (bounded by coverage), property-based (leveraging algebraic invariances), and shared-operator (through function reuse). This conceptual lens contextualizes our results and highlights where new architectural ideas are needed to achieve systematic compositionally. Overall, the coverage principle provides a unified lens for understanding compositional reasoning, and underscores the need for fundamental architectural or training innovations to achieve truly systematic compositionality. [^1] All code and resources are available at: https://github.com/kaistAI/coverage-principle
@@ 1 Introduction
Figure 1: Left: In a two-hop task $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, two fragments $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ satisfying $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ consistently yield the same final output when combined with the same context $x_{3}$, supporting their functional equivalence. Right: Among all possible inputs (few shown), we draw an edge between any two inputs that differ only by functionally equivalent fragments to form a substitution graph. Then, coverage is the set of observed inputs (highlighted as blue) and all inputs connected to them. The coverage principle states that a learner relying on pattern matching can reliably generalize only to inputs that are reachable from the observations by substituting functionally equivalent fragments.
Human generalization capability is often attributed to our ability to manipulate symbols compositionally [1, 2]. Despite the remarkable performance of Large Language Models (LLMs) [3, 4, 5, 6, 7, 8], our understanding of how they represent and manipulate compositional structures is still limited. Meanwhile, a growing body of work points to the brittleness of LLMs in performing systematic [^2] We refer to systematicity as an ability to generalize to unseen outcomes by exploiting their structural relations. For a more thorough discussion, see Fodor and Pylyshyn [9], Marcus [10], and Szabó [11]. compositional reasoning [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These observations suggest that LLMs succeed primarily through pattern matching, i.e., by exploiting surface-level statistical correlations between input fragments and outputs, rather than systematic compositional reasoning. However, prior analyses of the pattern-matching behavior remain largely observational and task-specific, leaving us without a unified framework to predict when they will succeed or fail.
We show that Transformers’ [26] generalization on synthetic compositional tasks is predictable through a unified framework grounded in one observation: Transformers generalize by utilizing functionally equivalent fragments supported by input-output relations. This insight allows us to move beyond observing success or failure cases to systematically predicting generalization boundaries. To this end, we first formalize the concept of functional equivalence, where two fragments (i.e., components of input sequences) are regarded as equivalent when they yield identical results in all shared contexts observed in the training data. Then, we introduce coverage, the set of inputs reachable from training data by substituting functionally equivalent fragments. Finally, we propose the coverage principle, which states that coverage is a necessary condition for compositional generalization, i.e., model predictions become unreliable beyond this fundamental boundary, as long as the only generalization mechanism is pattern matching.
From this conceptual formulation, we derive specific predictions about generalization in Transformers. First, we derive and empirically confirm that reliable generalization on a two-hop reasoning task requires training data to scale at least quadratically with the token set size. Remarkably, our experiments suggest that this scaling relationship remains invariant even for 20x larger model, underscoring that the limitation is inherent to data properties rather than model capacity.
Second, our framework also predicts specific complications in tasks with path ambiguities, where the same variable influences outputs through multiple computational paths. We show that Transformers cannot form unified representations of identical intermediate states in such cases, instead developing context-dependent state representations that fragment across different paths. This degrades both performance and interpretability, explaining why seemingly simple compositional tasks can remain challenging even for large models.
Third, while Chain-of-Thought (CoT) supervision [27, 28] improves data efficiency by ‘flattening’ tasks into sequences of single-hop problems, our analysis reveals a fundamental limitation: CoT-trained models still cannot resolve path ambiguities without showing nearly all possible in-domain combinations. This finding provides insights into understanding the Transformers’ failure modes in complex reasoning tasks with path ambiguities (e.g., planning [29, 30, 31]).
We place these findings within a broader, mechanism-based taxonomy that distinguishes three routes to generalization: (i) structure-based generalization via functional equivalence (the focus of this paper), (ii) property-based generalization that exploits algebraic invariances of individual primitive functions, and (iii) shared-operator generalization that reuses the same computation across positions. This taxonomy clarifies how certain model behaviours that occur outside the coverage region can arise from Type-II or Type-III mechanisms, providing a conceptual framework for analysing diverse sources of generalization.
Overall, the coverage principle offers a unified conceptual framework providing valuable insights into compositional generalization, highlighting how compositional reasoning and its mechanistic interpretation remain inherently bounded by coverage limitations. This underscores the need for architectural or training innovations to achieve truly systematic generalization in neural networks.
@@ 2 Related work
Compositional generalization and systematicity
Since Fodor and Pylyshyn [9] posed the systematicity challenge, arguing connectionist systems inherently lack symbolic structural capabilities, understanding how neural networks might achieve compositional reasoning has remained a persistent research question [32, 33, 34, 35, 36, 37, 38, 39, 10]. Central to this debate is the variable binding problem, how neural networks can flexibly associate variables with values across contexts [40, 41, 42], which Marcus [37] argues as the fundamental limitation of modern connectionist approaches. Despite AI advances, a growing body of work reveals limitations in neural networks’ compositional capabilities [14, 43, 18, 44, 45, 20, 46, 15, 21, 22, 23, 24, 47, 42]. Dziri et al. [17] observe that Transformers perform linearized subgraph matching to solve compositional tasks, but leave open when and why it occurs. Our coverage principle addresses this gap by formally demarcating what pattern matching models can achieve without true variable binding, showing both the possibilities and fundamental limitations of pattern matching-based approaches through data-based preconditions. This aligns with empirical findings of Andreas [48] and Auersperger and Pecina [49], where compositional data augmentation that implicitly expands coverage significantly improves generalization. Our derived power-law scaling relationship between token set size and required dataset size contributes to scaling laws literature [50, 51], but differs by characterizing a data requirement invariant to model size for compositional tasks. Unlike prior disentanglement work [52, 53, 54, 55] which focuss on latent factors, our data-centric view identifies fundamental boundary conditions for pattern-based generalization. The coverage principle thus complements recent architectural or training innovations [56, 57, 58, 59, 60] by showing what any pattern matching approach must overcome to achieve true systematicity.
Mechanistic interpretability
Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors [61, 62, 63, 64]. Recent work traces how Transformer components coordinate on compositional tasks [65, 66, 67, 68, 69]. Wang et al. [19] demonstrate that in-domain compositional generalization emerges through grokking, but do not identify necessary data conditions for such emergence. Our data-centric framework complements this optimization-centric view by pinpointing when generalizing circuits can exist: only when training data provides sufficient evidence of functional equivalence, essentially formalizing when the pattern matching mechanisms that Fodor and Pylyshyn [9] critique can succeed or must fail. Our findings also explain why standard interpretability techniques like logit lens [70, 71] may fail to identify compositional structure in models trained on tasks with path ambiguities.
3 A formulation of the coverage principle
@@
We now develop a formal framework of the coverage principle. We first provide an intuitive illustration with a two-hop compositional task, then generalize the intuitions to formalize the coverage principle for more general compositional tasks.
@@
Imagine a learner [^3] By “learner”, we mean any system that learns from data (e.g., neural networks). observing input-output pairs determined by a function $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$. The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is sampled from a finite set ${\mathcal{X}}$ [^4] For brevity, we use a shared token set ${\mathcal{X}}$. Position-specific domains ${\mathcal{X}}_{i}$ can be embedded as subsets of an enlarged token set $\tilde{{\mathcal{X}}}={\mathcal{X}}_{1}\cup{\mathcal{X}}_{2}\cup{\mathcal{X}}_{3}$ without loss of generality. Being unknown to the learner, let $f$ be a function that can be factorized as the composition of two primitive functions as $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig. 2(a). How can the learner generalize by only seeing the input-output pairs?
@@
Our key intuition is that a learner exploits the compositional structure only when observations show that two fragments behave identically. For instance, assume that $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ for $(x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{X}}^{2}$. From the underlying structure, these fragments are functionally equivalent, i.e., substituting one for the other never changes the output. If observations consistently support their equivalence, i.e., $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$ for observed $x_{3}$ values, this equivalence can be supported (Fig. 1 (Left)). Intuitively, the learner would be able to harness this equivalence to generalize to an unseen input $(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime\prime})$, provided the learner has seen $f(x_{1},x_{2},x_{3}^{\prime\prime})$.
@@
In short, the learner can utilize the observed functional equivalence to correctly infer the output of an unseen input, if it can reach an observed input by safe substitutions supported by observations (Fig. 1 (Right)). Coverage is a set of such inputs that are reachable from an observed input through chains of functionally equivalent substitutions. The essence of our framework, the coverage principle, is that a learner can only generalize inside the coverage when the only source of compositional generalization is the observation and utilization of functional equivalence.
@@
We now formalize these concepts for an arbitrary fixed-length task with an arbitrary set of observations. We restrict our attention to single-token prediction tasks defined as a deterministic mapping $f:{\mathcal{X}}^{n}\to{\mathcal{X}}$, where ${\mathcal{X}}$ is a finite set of tokens. We also consider a fixed observation set $D\subset{\mathcal{X}}^{n}$, a collection of inputs that are allowed to be observed by the learner. Write ${\bm{x}}=(x_{1},\dots,x_{n})\in{\mathcal{X}}^{n}$ and, for a subset $I\subset[n]:=\{1,\dots,n\}$, let ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ be a subsequence of ${\bm{x}}$. The first step is to formalize what it means for two subsequences to be functionally equivalent.
@@
Fix a nonempty proper subset of indices $I\subsetneq[n]$. Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{|I|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset D$ to be an $I$-co-occurrence of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ if it satisfies ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[n]\setminus I}={\bm{x}}^{\prime}_{[n]\setminus I}$. Also, we say ${\bm{a}}$ and ${\bm{a}}^{\prime}$ are functionally $k$-equivalent at $I$ in $D$, denoted by ${\bm{a}}\equiv^{I}_{D}{\bm{a}}^{\prime}$, if it satisfies:
@@
    (Sufficiency of co-occurrences.) There are $k$ or more distinct $I$-co-occurrences of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$;
@@
    (Consistency.) Every $I$-co-occurrence $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ satisfies $f({\bm{x}})=f({\bm{x}}^{\prime})$.
@@
Next, we ask: which inputs are reachable from observed data utilizing functional equivalence? To formalize this, we define the substitution graph: Let ${\mathcal{G}}_{D,k}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{n}$ of all possible inputs. Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[n]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$. With this substitution graph ${\mathcal{G}}_{D,k}$, we formally define the $k$-coverage as a set of inputs which are connected [^5] For an undirected graph ${\mathcal{G}}$, two vertices $u,v$ are connected if ${\mathcal{G}}$ contains a path between $u$ and $v$. to at least one observed input as follows:
@@
The $k$-coverage of $D$, denoted by $\mathrm{Cover}_{k}(D)$, is the set of all inputs in ${\mathcal{X}}^{n}$ that is connected to an input in $D$ on the substitution graph ${\mathcal{G}}_{D,k}$.
Note that the notion of coverage is a stricter condition of the conventional definition of in-domain, which is obtained by random train/test split [19] or taking combinations of observed internal computations [17]. We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms. Now, we formally phrase the coverage principle:
Claim 3.3 (Coverage principle).
Let a learner output $\hat{f}$ that is consistent with the training set $D$. If generalization is derived solely from observed functional $k$-equivalences and ${\bm{x}}\notin\mathrm{Cover}_{k}(D)$, the value $\hat{f}({\bm{x}})$ is unconstrained by $D$.
This establishes a fundamental boundary: no learner can reliably generalize beyond the coverage as long as the only source of the generalization is functional equivalence. In Sec. 5.1, we show that Transformers indeed fail to generalize beyond coverage when our random mapping dataset construction isolates functional equivalence as the sole generalization source. However, deeper insights come from asking when generalization success occurs in practice. While theory suggests that deep learning methods implicitly favor simpler solutions [72, 73, 74], which may involve leveraging functional equivalence, we still lack a systematic picture of how task structure, dataset, and model size interact to determine when such solutions emerge. We study this through experiments in Sec. 5.

### v2

# Characterizing Pattern Matching and Its Limits on Compositional Task Structures

@@ Abstract
Despite impressive capabilities, LLMs’ successes often rely on pattern-matching behaviors, yet these are also linked to OOD generalization failures in compositional tasks. However, behavioral studies commonly employ task setups that allow multiple generalization sources (e.g., algebraic invariances, structural repetition), obscuring a precise and testable account of how well LLMs perform generalization through pattern matching and their limitations. To address this ambiguity, we first formalize pattern matching as functional equivalence, i.e., identifying pairs of subsequences of inputs that consistently lead to identical results when the rest of the input is held constant. Then, we systematically study how decoder-only Transformer and Mamba behave in controlled tasks with compositional structures that isolate this mechanism. Our formalism yields predictive and quantitative insights: (1) Instance-wise success of pattern matching is well predicted by the number of contexts witnessing the relevant functional equivalence. (2) We prove a tight sample complexity bound of learning a two-hop structure by identifying the exponent of the data scaling law for perfect in-domain generalization. Our empirical results align with the theoretical prediction, under 20× parameter scaling and across architectures. (3) Path ambiguity is a structural barrier: when a variable influences the output via multiple paths, models fail to form unified intermediate state representations, impairing accuracy and interpretability. (4) Chain-of-Thought reduces data requirements yet does not resolve path ambiguity. Hence, we provide a predictive, falsifiable boundary for pattern matching and a foundational diagnostic for disentangling mixed generalization mechanisms.
@@ 1 Introduction
Despite the remarkable performance of Large Language Models (LLMs) (brown2020language; touvron2023llama), compositional generalization studies suggest that the core mechanism of generalization might be “pattern matching”, i.e., models learning local statistical regularities between input fragments and outputs in some cases (loula2018rearranging; johnson2017clevr; berglund2023reversal; wang2024grokked; mirzadeh2024gsm; keysers2019measuring; csordas2022ctl++). However, behavioral studies commonly employ task setups that allow multiple generalization sources (e.g., algebraic invariances, structural repetition), discussing pattern matching without a precise definition and diagnosing it post-hoc rather than characterizing it predictively. As a result, it remains unclear which behaviors should count as pattern matching and which should not, obscuring a constructive and testable account of its boundary.
To make this notion precise, we (1) introduce a model-agnostic, data-centric formalism for pattern matching and (2) systematically study how modern architectures like decoder-only Transformers (vaswani2017attention) and Mamba (gu2023mamba) perform generalization through pattern matching. Specifically, we first propose a model-agnostic and data-centric definition of pattern matching by formalizing the substitution of input patterns observed to result in identical outputs in shared contexts as functional-equivalence (Sec.˜3; henceforth, we use pattern matching as equivalent to functional-equivalence-based generalization). This induces a coverage boundary: if learning relies only on such evidence, reliable prediction is expected only for test inputs reachable by these substitutions.
Moreover, to isolate and study pattern-matching behaviors, we use controlled setups that deliberately remove other generalization sources and make functional equivalence the primary available mechanism. With this setting, our formalism yields predictive and quantitative insights about the limitations of pattern matching that, to our knowledge, have not been well characterized in prior works:
  • Generalization success is well predicted by the number of supporting contexts that witness the relevant functional equivalence. Mechanistically, Transformers implement functional equivalence via clustered intermediate representations at specific layers/positions, with clustering strength aligning with evidence strength (Sec.˜5).
  • We prove a tight sample complexity bound of pattern matching on a two-hop structure by identifying the exponent of the data scaling law (in terms of the token set size) which is necessary and sufficient for perfect in-domain generalization (Theorem˜6.1). The measured power-law exponent agrees with our theoretical bounds and remains stable under roughly $20\times$ parameter increase (from 68M to 1.5B) for GPT-2 (radford2019language), and also holds for Mamba (gu2023mamba) architecture (Sec.˜6).
  • When the same variable influences the output along multiple computational paths, models fail to form unified intermediate state representations. Analysis reveals that they instead develop context-dependent state representations, impairing both generalization and interpretability (Sec.˜7).
  • Chain-of-Thought (CoT) supervision (wei2022chain) reduces data requirements yet does not resolve path ambiguity without seeing nearly exhaustive in-domain combinations (Sec.˜8).
Finally, we situate this characterization of pattern matching within a mechanism-based taxonomy of generalization mechanisms, proposing two additional distinguishable mechanisms of generalization in compositional tasks: property-based and shared-operator generalization (Sec.˜9.2 and I).
Our formalism opens several research directions with practical implications (e.g., targeted data augmentation to maximize coverage) and motivates expansion to broader tasks and architectures, as well as systematic studies of how pattern matching interacts with other generalization mechanisms. Overall, our study provides a predictive, falsifiable boundary for what can be achieved through pattern matching alone and a foundational diagnostic for disentangling mixed mechanisms in modern neural networks.
@@
Figure 1: Illustration of functional equivalence. Left: In a two-hop task $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, two fragments $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ satisfying $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ consistently yield the same final output when combined with the same context $x_{3}$, supporting their functional equivalence. Right: Among all possible inputs (few shown), we draw an edge between any two inputs that differ only by functionally equivalent fragments to form a substitution graph. Then, coverage is the set of observed inputs (highlighted as blue) and all inputs connected to them. We define pattern matching as a type of generalization that occurs inside the coverage, harnessing functional equivalence.
@@ 2 Related Work
Pattern matching behaviors of LLMs on compositional tasks.   It is well perceived that pattern matching alone is inadequate for systematic generalization (fodor1988connectionism), and modern LLMs display generalization abilities that seem to be far beyond what pattern matching alone can do, as measured by their remarkable performance on complex benchmarks (achiam2023gpt). However, a growing body of work has consistently reported that LLMs still fall short on benchmarks designed to test compositionality (hupkes2020compositionality), including mathematical reasoning (mirzadeh2024gsm), multi-hop reasoning (yang2024large; wang2024grokked), and more (lake2018generalization; kim2020cogs; csordas2022ctl++; dziri2024faith). This gap between their capabilities and pattern-matching behaviors on compositional tasks calls for a principled framework to define what pattern matching is and to what extent a model’s behavior can be attributed to pattern matching, but it is mostly discussed with behavioral studies under the context of a specifically designed benchmark. Our work addresses this gap by formally defining pattern matching, and we systematically analyze models’ behaviors with controlled tasks that are designed to isolate pattern-matching regimes grounded in our framework.
Mechanistic interpretability.   Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors (elhage2021mathematical; olsson2022context; nanda2023progress; elhage2022toy). Recent work analyzes how Transformer components are causally related to certain behaviors (meng2022locating; hanna2023does; goldowsky2023localizing). In particular, it is reported that in-domain compositional generalization can emerge through grokking, with identifiable intermediate state representations inside Transformers (wang2024grokked). Our framework complements these works by providing mechanistic insights about pattern matching. Our findings also explain why standard interpretability techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) may fail to identify state representations in models trained on tasks with path ambiguities.
3 Formalizing Pattern Matching with Functional Equivalence
@@
We now develop a formal framework for pattern matching. We first provide an intuitive illustration with a two-hop structure, then generalize to arbitrary fixed-length discrete-sequence tasks.
@@
Imagine a learner observing data determined by $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$. The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is chosen from a finite set $\mathcal{X}$}. [^1] For brevity, we use a shared token set $\mathcal{X}$} in the main text. A more general notion using position-specific domains (e.g., ${\mathcal{X}}_{i}$’) can be used; see Sec. F.1. Suppose (unknown to the learner) that $f$ factorizes as the composition of two primitive functions, $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig.˜2(a). How can the learner generalize by only seeing the input-output patterns?
@@
Our key intuition is that a learner exploits the underlying patterns only when two fragments of inputs are observed to behave identically. For instance, assume that two fragments $(x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{X}}^{2}$ give the same implicit intermediate state upon the application of $f_{1}$, i.e., $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$. These fragments behave identically regardless of context, i.e., they are functionally equivalent: for all $x_{3}\in{\mathcal{X}}$, $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$. If observations consistently support their equivalence, i.e., $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$ for observed $x_{3}$ values, this equivalence can be supported (Fig.˜1 Left). Intuitively, the learner would harness this equivalence pattern to predict $f(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime\prime})$, provided the training set contains $f(x_{1},x_{2},x_{3}^{\prime\prime})$.
@@
Equivalently, the learner can utilize the observed functional equivalence to correctly infer the output of an unseen input, if it can reach an observed input by ‘safe substitutions’ (edges in the substitution graph) supported by observations (Fig.˜1 Right), which we define as a pattern matching. Coverage is a set of such inputs that are reachable from an observed input through chains of functionally equivalent substitutions. Then, coverage sets a boundary for what can be achieved by solely relying on substituting observed, equivalently behaving patterns. In other words, a learner can only generalize inside the coverage when it relies on functional equivalence, which we will define as pattern matching.
@@
We now formalize these concepts for an arbitrary fixed-length task with an arbitrary set of discrete sequence observations. We restrict our attention to single‑token prediction tasks defined as a deterministic mapping $f:{\mathcal{X}}^{\ell}\to{\mathcal{X}}$, where $\mathcal{X}$} is a finite set of tokens. We also consider a fixed observation set $D\subset{\mathcal{X}}^{\ell}$, a collection of inputs that are allowed to be observed by the learner. Write ${\bm{x}}=(x_{1},\dots,x_{\ell})\in{\mathcal{X}}^{\ell}$ and, for a subset $I\subset[\ell]:=\{1,\dots,\ell\}$, let ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ be a subsequence of $\bm{x}$}. The first step is to formalize what it means for two subsequences to be functionally equivalent.
@@
Fix a nonempty proper subset $I$ of indices in $\ell$]. Consider any set $S\subset{\mathcal{X}}^{\ell}$ of input sequences. [^2] The set $S$ can be any subset of the whole domain, e.g., $\mathcal{X}$$^{\ell}$ itself, the train dataset $D$, or whatever else. Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{\left|I\right|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ to be an $I$-co-occurrence of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$ if it satisfies $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset S$ and ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[\ell]\setminus I}={\bm{x}}^{\prime}_{[\ell]\setminus I}$. Also, the subsequences $\bm{a}$} and $\bm{a}$$^{\prime}$ are said to be functionally $k$-equivalent at $I$ in $S$ and denoted by ${\bm{a}}\equiv^{I}_{S}{\bm{a}}^{\prime}$, if it satisfies:
@@
    (Sufficiency of co-occurrences.) There are $k$ or more distinct $I$-co-occurrences of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$;
@@
    (Consistency.) Every $I$-co-occurrence $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$ satisfies $f({\bm{x}})=f({\bm{x}}^{\prime})$.
@@
Next, we ask: which inputs are reachable from observed data utilizing functional equivalence? To formalize this, we define substitution graph: Let ${\mathcal{G}}^{(D,k)}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{\ell}$ of all possible inputs. Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[\ell]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$. This process is illustrated on the right side of Fig.˜1, as a special case where $k=1$. With this substitution graph ${\mathcal{G}}^{(D,k)}$, we formally define the $k$-coverage as a set of inputs which are connected [^3] For an undirected graph $\mathcal{G}$}, two vertices $u,v$ are connected if $\mathcal{G}$} contains a path between $u$ and $v$. to at least one observed input as follows:
@@
The $k$-coverage of $D\subset{\mathcal{X}}^{\ell}$, denoted by $\mathrm{Cover}_{k}(D)$, is the set of all inputs in $\mathcal{X}$$^{\ell}$ connected to an ${\bm{x}}\in D$ with a path in the substitution graph ${\mathcal{G}}^{(D,k)}$.
Note that the notion of coverage is a stricter condition of the canonical definition of in-domain (ID), which is obtained by random train/test split (wang2024grokked) or taking combinations of observed internal computations (dziri2024faith). In Sec.˜5, we demonstrate that learners may not necessarily generalize on data that are classified as ID in a canonical sense, but coverage can precisely explain when and why this occurs. We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms, and we demonstrate that the predictions made by our framework are invariant across model architecture and scale in Sec.˜6 and 7. Finally, $k$-coverage can be algorithmically determined for any fixed-length discrete sequence tasks (Alg. 1), which we use for the analyses in the following sections.
Now, we formally define pattern matching as a kind of generalization that is done by substituting functionally $k$-equivalent fragments of inputs, whose boundary is precisely the $k$-coverage defined above. This formalization enables us to predict, before testing, which inputs will be reliably handled through pattern matching and which require additional mechanisms. In other words, we view pattern matching as possible only within $k$-coverage, and generalization outside the coverage requires generalization mechanisms other than pattern matching, which we discuss in Sec.˜10 and I. In the following sections, we draw a systematic picture of how task structure, dataset, and model size interact to determine the success and failure of pattern matching through controlled setups, leading us to important and nontrivial insights.